Search Previous Year Questions

Find real questions from previous competitive exam papers with answers and solutions.

Showing 10 question(s)
Clear

Question 111

National Defence Academy (NDA) 2025• Mathematics

A fair coin is tossed 4 times. What is the probability that two heads do not occur consecutively?

  1. AOption (a)

    1/8

  2. BOption (b)

    3/8

  3. COption (c)

    7/16

  4. DOption (d)

    1/2

View answer & explanation

Correct answer

Answer: (d)D.

1/2

Why this answer is correct

Calculation:

Total number of outcomes for 4 coin tosses:

n(S)=24=16 n(S) = 2^4 = 16

Favorable outcomes with two consecutive heads:

⇒ E=HHTT, THHT, TTHH, HHHT, HTHH, HHTH, THHH, HHHH E = { \text{HHTT, THHT, TTHH, HHHT, HTHH, HHTH, THHH, HHHH} }

⇒ n(E)=8⇒n(E′)=16−8=8 n(E) = 8 \Rightarrow n(E') = 16 - 8 = 8

Probability of not getting two consecutive heads:

⇒ P(E′)=816=12 P(E') = \frac{8}{16} = \frac{1}{2}

Hence, the correct answer is Option 4.

Question 112

National Defence Academy (NDA) 2025• Mathematics

In a throw of three dice, what is the probability of getting one prime number, one composite number and one number which is neither prime nor composite?

  1. AOption (a)

    1/2

  2. BOption (b)

    1/3

  3. COption (c)

    1/4

  4. DOption (d)

     1/6

View answer & explanation

Correct answer

Answer: (d)D.

 1/6

Why this answer is correct

Calculation:

Given:

Prime numbers = {2, 3, 5} ⇒ 3 choices

Composite numbers = {4, 6} ⇒ 2 choices

Neither prime nor composite = {1} ⇒ 1 choice

Probability of selecting one prime, one composite, and one neither prime nor composite:

P=36⋅26⋅16⋅3! P = \frac{3}{6} \cdot \frac{2}{6} \cdot \frac{1}{6} \cdot 3!

Calculating:

=6216⋅6=36216=16 = \frac{6}{216} \cdot 6 = \frac{36}{216} = \frac{1}{6}

Hence, the correct answer is Option 4.

Question 113

National Defence Academy (NDA) 2025• Mathematics

An integer is chosen at random from the first 50 integers. What is the probability that the integer is neither divisible by 5 nor 9?

  1. AOption (a)

     7/10

  2. BOption (b)

    18/25

  3. COption (c)

    37/50

  4. DOption (d)

    19/25

View answer & explanation

Correct answer

Answer: (b)B.

18/25

Why this answer is correct

Calculation:

Given:

Numbers divisible by 5 = 10

Numbers divisible by 9 = 5

Numbers divisible by both 5 and 9 = 1

Using the inclusion-exclusion principle:

Divisible by 5 or 9 =10+5−1=14 = 10 + 5 - 1 = 14

Total numbers = 50

Not divisible by 5 or 9 =50−14=36 = 50 - 14 = 36

Required probability:

P=3650=1825 P = \frac{36}{50} = \frac{18}{25}

Hence, the correct answer is Option 2.

Question 114

National Defence Academy (NDA) 2025• Mathematics

Out of 50 consecutive natural numbers, two integers are chosen at random. What is the probability that their sum is odd?

  1. AOption (a)

    1/2

  2. BOption (b)

    24/49

  3. COption (c)

    1/4

  4. DOption (d)

    25/49

View answer & explanation

Correct answer

Answer: (d)D.

25/49

Why this answer is correct

Calculation:

Out of 50 consecutive natural numbers:

Even numbers = 25

Odd numbers = 25

Sum of one even and one odd number is always odd.

Number of ways to choose one even and one odd:

=25⋅25=625 = 25 \cdot 25 = 625

Total number of ways to choose any two numbers from 50:

=50C2=50⋅492=1225 = {}^{50}C_2 = \frac{50 \cdot 49}{2} = 1225

Required probability:

P(sum is odd)=6251225=2549 P(\text{sum is odd}) = \frac{625}{1225} = \frac{25}{49}

Hence, the Correct answer is Option 4.

Question 115

National Defence Academy (NDA) 2025• Mathematics

The standard deviation of 100 observations is 10. If 20 is added to each observation, then what will be the new standard deviation?

  1. AOption (a)

    10

  2. BOption (b)

    15

  3. COption (c)

    20

  4. DOption (d)

    25

View answer & explanation

Correct answer

Answer: (a)A.

10

Why this answer is correct

Calculation:

Given:

Standard deviation of 100 observations = 10

A constant value 2020 is added to each observation.

Property of standard deviation:

Adding or subtracting a constant from each observation does not change the standard deviation

Hence,

New standard deviation = Old standard deviation = 10

Hence, the correct answer is Option 1.

Question 116

National Defence Academy (NDA) 2025• Mathematics

Let X be a random variable following binomial distribution with parameters n = 5 and p = k Further, P(X = 1) = 0.4096 and P(X = 2) = 0.2048 . What is the value of k? 

  1. AOption (a)

    0.2

  2. BOption (b)

    0.25

  3. COption (c)

     0.3

  4. DOption (d)

     0.35

View answer & explanation

Correct answer

Answer: (a)A.

0.2

Why this answer is correct

Calculation:

Given:

n=5 n = 5

p=k p = k , q=1−k q = 1 - k

Binomial probabilities:

⇒ P(X=1)=5C1⋅k⋅(1−k)4=0.4096 P(X = 1) = {}^5C_1 \cdot k \cdot (1 - k)^4 = 0.4096

⇒ P(X=2)=5C2⋅k2⋅(1−k)3=0.2048 P(X = 2) = {}^5C_2 \cdot k^2 \cdot (1 - k)^3 = 0.2048

Taking ratio:

⇒ 5C2⋅k2⋅(1−k)35C1⋅k⋅(1−k)4=0.20480.4096 \frac{{}^5C_2 \cdot k^2 \cdot (1 - k)^3}{{}^5C_1 \cdot k \cdot (1 - k)^4} = \frac{0.2048}{0.4096}

Simplifying:

⇒ 10⋅k5⋅(1−k)=12⇒4k=1−k⇒5k=1⇒k=15=0.2 \frac{10 \cdot k}{5 \cdot (1 - k)} = \frac{1}{2} \Rightarrow 4k = 1 - k \Rightarrow 5k = 1 \Rightarrow k = \frac{1}{5} = 0.2

Hence, the correct answer is Option 1.

Question 117

National Defence Academy (NDA) 2025• Mathematics

The frequency distribution of the marks obtained by students in a Science examination is given below:

 

Marks5-15
15-25
25-35
35-45
Number of
students
20303020

What is the arithmetic mean?

  1. AOption (a)

    20

  2. BOption (b)

    25

  3. COption (c)

    30

  4. DOption (d)

    35

View answer & explanation

Correct answer

Answer: (b)B.

25

Why this answer is correct

Calculation:

The given data is a grouped frequency distribution.

 

Class Intervalfixixifi
5 – 152010200
15 – 253020600
25 – 353030900
35 – 452040800
Total∑fi=1002500

Compute the totals:

⇒ ∑fi=20+30+30+20=100 \sum f_i = 20+30+30+20 = 100

⇒ ∑xifi=200+600+900+800=2500 \sum x_i f_i = 200+600+900+800 = 2500

Mean of the distribution is given by:

⇒ Mean=∑xifi∑fi \text{Mean} = \frac{\sum x_i f_i}{\sum f_i}

Substituting the values:

⇒ Mean=2500100=25 \text{Mean} = \frac{2500}{100} = 25

Hence, the correct answer is Option 2.

Question 118

National Defence Academy (NDA) 2025• Mathematics

If P(A) = 0.3 , P(B) = 0.4 and P( A |B)=0.5, then what is the value P( B |A) ?

  1. AOption (a)

    0.325

  2. BOption (b)

     0.333

  3. COption (c)

    0.375

  4. DOption (d)

    0.667

View answer & explanation

Correct answer

Answer: (d)D.

0.667

Why this answer is correct

Calculation:

Given:

 P(A) = 0.3 , P(B) = 0.4 and P( A |B)=0.5

Using the formula of conditional probability:

P!(AB)=P(A∩B)P(B) P!\left(\frac{A}{B}\right)=\frac{P(A\cap B)}{P(B)}

Substitute the given values:

0.5=P(A∩B)0.4 0.5=\frac{P(A\cap B)}{0.4}

Solving for P(A∩B)P(A\cap B):

P(A∩B)=0.5×0.4=0.2 P(A\cap B)=0.5\times 0.4=0.2

Now find P!(BA)P!\left(\frac{B}{A}\right):

P!(BA)=P(A∩B)P(A) P!\left(\frac{B}{A}\right)=\frac{P(A\cap B)}{P(A)}

Substitute the values:

P!(BA)=0.20.3=23≈0.667 P!\left(\frac{B}{A}\right)=\frac{0.2}{0.3}=\frac{2}{3}\approx 0.667

Hence, the correct answer is Option 4.

Question 119

National Defence Academy (NDA) 2025• Mathematics

If P(A) = 1/3 , P(B) = 1/2 and P(A∩B)=1/4 P(A\cap B)=1/4 , then what is the value of P(A‾∪B)P(\overline{A}\cup B) ?

  1. AOption (a)

    ​7/12

  2. BOption (b)

     2/3

  3. COption (c)

    3/4

  4. DOption (d)

    11/12

View answer & explanation

Correct answer

Answer: (d)D.

11/12

Why this answer is correct

Calculation:

Given:

P(A)=13,P(B)=12,P(A∩B)=14 P(A)=\frac{1}{3},\quad P(B)=\frac{1}{2},\quad P(A\cap B)=\frac{1}{4}

We know that:

⇒ P(Aˉ)=1−P(A) P(\bar{A})=1-P(A)

Using the formula:

⇒ P(Aˉ∪B)=P(Aˉ)+P(B)−P(Aˉ∩B) P(\bar{A}\cup B)=P(\bar{A})+P(B)-P(\bar{A}\cap B)

Substitute P(Aˉ)=1−P(A)P(\bar{A})=1-P(A):

⇒ P(Aˉ∪B)=1−P(A)+P(B)−[P(B)−P(A∩B)] P(\bar{A}\cup B)=1-P(A)+P(B)-[P(B)-P(A\cap B)]

Substitute the given values:

=1−13+12−(12−14) =1-\frac{1}{3}+\frac{1}{2}-\left(\frac{1}{2}-\frac{1}{4}\right)

Simplifying:

=23+12−14 =\frac{2}{3}+\frac{1}{2}-\frac{1}{4}

=8+6−312=1112 =\frac{8+6-3}{12}=\frac{11}{12}

Hence, the correct answer is Option 4.

Question 120

National Defence Academy (NDA) 2025• Mathematics

Consider the following statements:

I. Mean and variance have the same unit of measurement.

II. Mean deviation and standard deviation have the same unit of measurement.

Which of the statements given above is/are correct?

  1. AOption (a)

    I only

  2. BOption (b)

    II only

  3. COption (c)

    Both I and II

  4. DOption (d)

     Neither I nor II

View answer & explanation

Correct answer

Answer: (b)B.

II only

Why this answer is correct

Calculation:

Statement I: Mean and variance have the same unit of measurement.

  • Mean has the same unit as the original data.
  • Variance is the square of the unit of the data (since deviations are squared).
  • Therefore, mean and variance do NOT have the same unit.
  • Statement I is incorrect.

Statement II: Mean deviation and standard deviation have the same unit of measurement.

  • Mean deviation is based on absolute deviations, so it has the same unit as the data.
  • Standard deviation is the square root of variance, hence it also has the same unit as the data.
  • Therefore, Statement II is correct.

Hence, the correct answer is Option 2.